# Entangling gates from cabling of knots

### v1

@@ Abstract
While there is a general consensus about the structure of one qubit operations in topological quantum computer, two qubits are as usual a more difficult and complex story of different attempts with varying approaches, problems and effectiveness. In this paper we discuss how to construct an efficient realization of a two qubit gate in topological quantum computer, by using principle of cabling from the knot theory. This allows to construct a braiding of cables dependent on the parameters of the theory where there is a low probability of moving out of computational space (high fidelity of operation) while there is a non-trivial entangling two-qubit operation. We also present some examples of these operations for different parameters of the theory.
@@ 1 Introduction
States of anyons and anti-anyons correspond to particular states of qubits, while intertwining of their world-lines in a three-dimensional space can be seen as operations as they change the states of the particles. Explicitly it means that operations are quantum $\mathcal{R}$-matrices of $U_{q}(SU(N))$ [11].
One of the possibilities to define the computational space of the qubit, as was extensively discussed in [9, 10, 12, 13, 14, 15], is to consider plat representation of knot or link where state of the qubit corresponds to the space of irreducible representations in two anyon-anti-anyon pairs (see Fig.1 for an example). We discuss the explicit construction in the next section.
@@ 2 One-qubit operations
The initial state that is depicted in a lower block of Fig.1 corresponds to the creation of two anyon pairs. Upon creation two anyons in each pair are in conjugate representations and pair as whole is projected onto trivial representation. After intertwining, which corresponds to an action of $\mathcal{R}$-matrices, the representation of a pair can change. Still, braid as a whole remains in trivial representation, since $\mathcal{R}$-matrices do not change the irreducible representations travelling along any number of strands. We assume for simplicity, that each strand is in fundamental/anti-fundamental representation, but in a particular experimental realization it is fixed by the types of anyons we would be able to create. As it was previously suggested, [9, 10, 12, 13, 14, 15] the two-bridge knot, i.e. knot with two initial pairs of anyons, corresponds to one qubit computing. Indeed, the projection onto trivial representation of the product of two fundamental and two anti-fundamental representations is two-dimensional. To get the trivial representation in the whole braid, each of the two pairs of strands should be either in trivial or adjoint representation. It can be seen, for example, as follows.
@@
Vectors $|V_{\emptyset}\rangle$ and $|V_{1}\rangle$ form a Hilbert space of one qubit. Let us stress, that computational space of a qubit is the space of different irreducible representations rather than an internal vector space of a particular irreducible representation. The closest analogy are $SU(3)$ and isospin multiplets of particles. Anyons in different representations correspond to particles in different hypermultiplets, in other words, different particles.
@@
where $q=\text{exp}(\frac{2\pi i}{k+N})$ and $A=q^{N}$ with $k$ and $N$ being parameters of Chern-Simons theory (1). These matrices correspond to the intersections of two nearby strands ($T$-operators) and to the change of basis from intersection of $1$ and $2$ strands to $2$ and $3$ ($S$-operators). As was discussed in [12, 13] these operators form a universal set of one-qubit gates (universality holds for qudits upon generalization to higher representations).
Thereby, after two pairs of anyon and anti-anyon are created (the initial state of the quantum computer), operations on a computational space are performed with operators (3) in accordance with particular braid: each intersection of two strands corresponds to a multiplication of a particular operator. Last part is the annihilation of anyon-anti-anyon pairs [^1] Note that due to intersections, the annihilated particles in each pair might differ from initial ones which corresponds to the projection onto trivial representation of each pair of strands. The process of annihilation has non-trivial probability, which is the result of a quantum program. From the representations point of view, braid (algorithm) corresponds to a two-by-two matrix $\mathcal{B}_{XY}$, which is a product of matrices from (3) [^2] Let us remind that $X$ and $Y$ takes values in space of different irreducible representations. While the full program with initial creation and measurement corresponds to a matrix element $\mathcal{B}_{\emptyset\emptyset}$, it describes the probability amplitude of two pairs of anyons being created, then entangled and then annihilated.
@@ 3 Two-qubits and the entangling operation
Figure 2: Operation, which entangles two qubits, can be constructed by braiding cables of two strands to form a more complex knot. If we entangle them by braiding individual strands the hilbert space of the structure will be much larger than the computational space of two qubits.
@@
In full analogy to the case of a four strands diagrams, now we should take a product of representations on all eight strands and then take all trivial representations which appear there. In the case of $U_{q}(SU(N))$ group we take the following product:
@@
After this type of entangling, (see for example Fig.3), representations transform in the following way. Trivial representation $\emptyset$ interacts trivially with all other representations. Therefore, element $I.$ will not be changed by any braiding. Elements $II.$ and $III.$ conserve representations too, as long as cables do not change the position, i.e. one keeps the left pair on the left in the end of the intertwining and the right pair on the right [^5] It is possible also to move the right pair as a whole to the left, and the left pair to the right, but this is more difficult to understand and study the corresponding operator. However, if we separate two cables and move only one of them to the other side, the whole structure will be destroyed.. Since two of the cables on one side are in trivial representation and do not interact with other cables, there is basically no two-qubit interaction. The whole braid can be reduced to the one-qubit interactions on the side with two adjoint representations (right side in case $II.$ or left side in case $III.$).
@@ 4 $U_{q}(SU(2))$ braiding
Due to the properties of the $U_{q}(SU(2))$ algebra all the representations are self-conjugate. Therefore, the matrices $S$ and $\bar{S}$ coincide. Similarly matrices $T$ and $\bar{T}$ should also coincide, but up to a general coefficient. This is due to the different additional property of these matrices. There are in fact three different braiding matrices. Depending on a set of intersections 4-plat picture can correspond either to a knot or to a link. In the latter case there are two separate link components, each with its own representation. While the representation and its conjugate are equivalent in the $U_{q}(SU(2))$ algebra, they are not exactly equal in the same basis. This means that the system should not be able to distinguish them when the strands correspond to two different components, because the basis is not fixed between the different components. We will denote this crossing by $T_{i}$, and the operator is given by the formulae in [24]. However, the matrices that correspond to the intersection of the one same component can depend on the direction of the strands, or, in other words they can distinguish representation and its conjugate. These matrices are defined by the properties on Fig.4 and are equal to:
@@
Using these matrices one can construct an operator $\mathcal{B}$ that corresponds to the braiding of the adjoint strands. We are interested in the matrix element $\mathcal{B}_{1,1}$ of this matrix, which corresponds to the transformation of two-qubits that are initially in a computational space (part $IV.$ from (7)) and are kept there in the end. Therefore, probability of two qubits to remain in computable space $\mathcal{P}$ and phase shift $\phi$ are given by

### v2

@@ Abstract
While there is a general consensus about the structure of one qubit operations in topological quantum computer, two qubits are as usual a more difficult and complex story of different attempts with varying approaches, problems and effectiveness. In this paper we discuss how to construct an efficient realization of a two qubit gate in topological quantum computer, by using principle of cabling from the knot theory. This allows oneto construct a braiding of cables dependent on the parameters of the theory where there is a low probability of moving out of computational space (high fidelity of operation) with a non-trivial entangling two-qubit operation. We also present some examples of these operations for different parameters of the theory.
@@ 1 Introduction
States of anyons and anti-anyons correspond to particular states of qubits, while intertwining of their world-lines in a three-dimensional space can be seen as operations as they change the states of the particles.
Observables in quantum Chern-Simons theory possess a more complex symmetry than a $SU(N)$ we put in the action. Namely they display properties, characteristic to the quantum group (quantum algebra) $U_{q}(SU(N))$, [9]. It is a deformation of universal enveloping algebra of the $SU(N)$ and is widely studied in relation to the different models in theoretical and mathematical physics. In the context of this paper it is important that the crossings of the world lines correspond to the quantum $\mathcal{R}$-matrices of $U_{q}(SU(N))$ [10].
One of the possibilities to define the computational space of the qubit, as was extensively discussed in [11, 12, 13, 14, 15, 16], is to consider 4-plat representation of knot or link where state of the qubit corresponds to the space of irreducible representations in two anyon-anti-anyon pairs (see Fig.1 for an example). We discuss the explicit construction in the next section.
@@ 2 One-qubit operations
The initial state, that is depicted in a lower block of Fig.1 corresponds to the creation of two anyon pairs. Upon creation two anyons in each pair are in conjugate representations, and pair as a whole is projected onto trivial representation. After intertwining, which corresponds to an action of $\mathcal{R}$-matrices, the representation of a pair can change. Still, braid as a whole remains in trivial representation, since $\mathcal{R}$-matrices do not change the irreducible representations travelling along any number of strands. We assume for simplicity, that each strand is in fundamental/anti-fundamental representation, however a particular experimental realization may be built on anyons in different representations which would depend on the types of anyons we would be able to create. As it was previously suggested, [11, 12, 13, 14, 15, 16] the 4-plat knot (also known as two-bridge knots), i.e. knot with two initial pairs of anyons, corresponds to one qubit computing. Indeed, the projection onto trivial representation of the product of two fundamental and two anti-fundamental representations is two-dimensional. To get the trivial representation in the whole braid, each of the two pairs of strands should be either in trivial or adjoint representation. It can be seen, for example, as follows:
@@
Vectors $|V_{\emptyset}\rangle$ and $|V_{1}\rangle$ form a Hilbert space of one qubit. Let us stress, that computational space of a qubit is the space of different irreducible representations rather than an internal vector space of a particular irreducible representation. The closest analogy are $SU(3)$ and isospin multiplets of particles. Anyons in different representations correspond to particles in different hypermultiplets, in other words, to different particles.
@@
where $q=\text{exp}(\frac{2\pi i}{k+N})$ and $A=q^{N}$ with $k$ and $N$ being parameters of Chern-Simons theory (1). In the $SU(N)$ case anyons and anti-anyons are different particles and therefore operators for the crossings of the trajectories depend on the exact type of particles which are crossing. If they are of the same type then operator $T$ should be chosen, if they are different then the operator is $\bar{T}$. $T$ and $\bar{T}$ are diagonalized $\mathcal{R}$-matrices written for the particular representations. Also since we write operators which act on all four anyons, we cannot diagonalize all of the at the same time. Matrices $T$ and $\bar{T}$ are diagonalized operators, while operators $S$ and $\bar{S}$ describe the change of bases to make different crossing operators diagonal. They also depend on the placement of particles of different types therefore there are two of them. See [13, 14] for more details. As was discussed in [13, 14] these operators form a universal set of one-qubit gates (universality holds for qudits upon generalization to higher representations).
Thereby, after two pairs of anyon and anti-anyon are created (the initial state of the quantum computer), operations on a computational space are performed with operators (3) in accordance with particular braid: one should take a product of operators corresponding to the different crossings in an order defined by a braid. Last part is the annihilation of anyon-anti-anyon pairs [^1] Note that due to intersections, the annihilated particles in each pair might differ from initial ones. which corresponds to the projection onto trivial representation of each pair of strands. The process of annihilation has non-trivial probability, which is the result of a quantum program. From the representations point of view, braid (algorithm) corresponds to a two-by-two matrix [^2] Let us remind that $X$ and $Y$ takes values in space of different irreducible representations. $\mathcal{B}_{XY}$, which is a product of matrices from (3). The full program with initial creation and measurement corresponds to a matrix element $\mathcal{B}_{11}$, when $X=Y=\emptyset$, it describes the probability amplitude of two pairs of anyons being created, then entangled and then annihilated.
@@ 3 Two-qubits and the entangling operation
Figure 2: Operation, which entangles two qubits, can be constructed by braiding cables of two strands to form a more complex knot. If we entangle them by braiding individual strands, the Hilbert space of the structure will be much larger than the computational space of two qubits.
@@
In full analogy to the case of a four strands diagrams, now we should take a product of representations on all eight strands and then take all trivial representations, which appear there. In the case of $U_{q}(SU(N))$ group we take the following product:
@@
There is no case which includes $\emptyset\otimes\text{adj}$ on one side since we want to start from the computable space where both sides are in trivial representation, which cannot be achieved if there is a product $\emptyset\otimes\text{adj}$.
After this type of entangling, (see for example Fig.3), representations transform in the following way. Trivial representation $\emptyset$ interacts trivially with all other representations. Therefore, case $I.$ will not be changed by any braiding. Cases $II.$ and $III.$ also do not change as long as cables do not change the position, i.e. one keeps the left pair on the left in the end of the intertwining and the right pair on the right [^5] It is possible also to move the right pair as a whole to the left, and the left pair to the right, but this is more difficult to understand and study the corresponding operator. However, if we separate two cables and move only one of them to the other side, the whole structure will be destroyed.. Since two of the cables on one side are in trivial representation and do not interact with other cables, there is basically no two-qubit interaction. The whole braid can be reduced to the one-qubit interactions on the side with two adjoint representations (right side in case $II.$ or left side in case $III.$).
@@ 4 $U_{q}(SU(2))$ braiding
Due to the properties of the $U_{q}(SU(2))$ algebra, all the representations are self-conjugate. Therefore, the matrices $S$ and $\bar{S}$ coincide. Similarly matrices $T$ and $\bar{T}$ should also coincide, but up to a general coefficient. This is due to the different additional property of these matrices. There are in fact three different braiding matrices. Depending on a set of intersections, 4-plat picture can correspond either to a knot or to a link. In the latter case there are two separate link components, each with its own representation. While the representation and its conjugate are equivalent in the $U_{q}(SU(2))$ algebra, they are not exactly equal in the same basis. This means that the system should not be able to distinguish them when the strands correspond to two different components, because the basis is not fixed between the different components. We will denote this crossing by $T_{i}$, and the operator is given by the formulae in [9]. However, the matrices that correspond to the intersection of the one same component can depend on the direction of the strands, or, in other words they can distinguish representation and its conjugate. These matrices are defined by the properties on Fig.4 and are equal to:
@@
Using these matrices one can construct an operator $\mathcal{B}$ that corresponds to the braiding of the adjoint strands. We are interested in the matrix element $\mathcal{B}_{1,1}$ of this matrix, which corresponds to the transformation of two-qubits that are initially in a computational space (case $IV.$ from (7)) and are kept there in the end. Therefore, probability of two qubits to remain in computable space $\mathcal{P}$ and phase shift $\phi$ are given by